OpenStudy (anonymous):

Solve by completing the square: 3x^2+12x+14=0

OpenStudy (anonymous):

3x^2 + 12x + 14 = 0 3(x^2 + 4x + 14/3) = 0 x^2 + 4x + 14/3 = 0 x^2 + 4x = -14/3 x^2 + 4x + (4/2)^2 = -14/3 + (4/2)^2 x^2 + 4x + 4 = -14/3 + 4 (x + 2)^2 = 4 - 14/3 x + 2 = ±sqrt(4 - 14/3) x = -2 ±sqrt(4 - 14/3)